# Moran Mizrahi1 Chen Shani2 Gabriel Stanovsky1 Dan Jurafsky2 Dafna Shahaf1 1

### v1

@@ 6.1 Experimental Details
Here we describe the process of selecting prompting parameters for GPT-4o as well as the human evaluation setup, both of which were consistently applied across both experiments.
Prompt and Temperature Selection. LLMs are known to be sensitive to prompt paraphrases Sclar et al. (2023); Mizrahi et al. (2024); Voronov et al. (2024). To optimize our parameters, we first determined the temperature value by generating outputs from three random prompts. After testing various settings, we selected $t=1$ as the highest temperature that produced coherent recipes, noting that higher values resulted in gibberish. With the temperature fixed, and following the methodology of Mizrahi et al. (2024), a team member curated 100 prompts per experiment, generating three recipes for each prompt. The 300 resulting recipes were then reviewed, and the more unique outputs were tagged. Ultimately, the final prompt was chosen for its ability to consistently generate creative recipes across all outputs (see the chosen prompts for both experiments in Appendix G).
@@
Human Evaluation Setup. We used Prolific to recruit and manage a total of 48 participants for both experiments. We pre-screened participants based on their cooking frequency, comfort with adjusting recipes, and ability to judge creative outcomes. Participants were paid an estimated hourly rate of £9, adhering to ethical compensation standards. Participants were asked to rate random recipes drawn from the two models on various value and novelty aspects (see the list of question in Appendix H). To reduce cognitive load, they viewed short recipe summaries (also generated using GPT-4o), with the option to view full recipes if desired. To mitigate ranking noise, each recipe was rated by five annotators, and final scores were determined using median ranks. The novelty and value scores were set to be the means of their corresponding questions, and the value score was also binarized using a threshold of 4 on three related value questions. On average, each participant rated 31.25 recipes (std=33.827). To ensure quality, we excluded participants who, on average, completed reading a recipe and answering its eight questions in less than 45 seconds.
Figure 5: Ingredient frequencies in the original recipe repository (in pink) and the models’ generated recipes. The black histogram shows $\mathcal{D}$ishCover closely follows the repository’s distribution. GPT-4o (blue) shows spikes, highlighting bias to certain ingredients.
@@ 6.2 Diversity in Generated Recipes
This difference was also reflected in the average ingredient counts per recipe: GPT-4o’s recipes had nearly twice as many ingredients as $\mathcal{D}$ishCover’s: in the first experiment, $\mathcal{D}$ishCover used 12.3 ingredients on average (std=3.71) compared to GPT-4o’s 24.98 (std=5.2); in the second experiment, $\mathcal{D}$ishCover used 13.32 ingredients on average (std=4.39) vs. GPT-4o’s 23.79 (std=3.25). For reference, the average ingredient count in the recipe repository was lower, 9.33 (std=4.31). This imbalance suggests that GPT-4o tends to generate separate sub-dishes, each with its own ingredients, and then combine them into a final dish rather than truly blending the culinary ideas.
Figure 6: t-SNE visualization for recipe embeddings, experiment 1. Shapes denote dish pairs, colors denote models. The blue clusters (GPT-4o) of the same shape appear tight and localized, while the pink ($\mathcal{D}$ishCover) are scattered, showing higher diversity across $\mathcal{D}$ishCover’s outputs.
Figure 7: t-SNE visualization for recipe embeddings, experiment 2. Gray points represent 1K random recipes from the general repository. Again, $\mathcal{D}$ishCover’s outputs are more diverse.
@@
GPT-4o showed a strong fixation on certain structures and ingredients. When asked to merge two dishes, it often followed the same approach in multiple attempts. For instance, when asked to fuse “lentil soup” and “jam”, GPT-4o repeatedly generated variations of lentil soup served alongside a separately prepared jam, typically plated in the same way (e.g., placing a dollop of jam in the center of the soup). This fixation persisted when asked to combine specific given recipes for each dish, and even when asked to merge broader dish types (e.g., a general soup with a general dip).
Similarly, GPT-4o repeatedly used the same (uncommon) ingredients. For instance, 56% of its recipes included smoked paprika (compared to just 0.375% in the repository), 47% used mint, 41% maple syrup, and 29% pomegranate molasses. Figure 5 shows ingredient frequencies in GPT-4o and $\mathcal{D}$ishCover recipes against the repository baseline. While $\mathcal{D}$ishCover’s distribution aligns closely with the original data, GPT-4o has a strong bias toward certain low-frequency ingredients.
@@
Quantifying Diversity via Embeddings.
We also analyzed the embeddings of the generated recipes using a fine-tuned Sentence-BERT model specialized for cooking recipes (see Appendix A). Figure 6 presents a t-SNE visualization for the first experiment, where each shape represents a dish pair, and color indicates the model ($\mathcal{D}$ishCover in pink, GPT-4o in blue). GPT-4o’s recipes formed tight clusters for each pair, whereas $\mathcal{D}$ishCover’s were more dispersed. A similar pattern emerges in experiment 2 (Figure 7). By construction, $\mathcal{D}$ishCover’s embeddings are widely distributed. The interesting thing to notice, however, is that GPT-4o’s embeddings tend to cluster closely. Figure 8 reinforces these findings with heatmaps of cosine similarities among all recipe embeddings. In the first experiment, $\mathcal{D}$ishCover’s outputs had an average similarity of 0.387 (std=0.120), while GPT-4o’s were higher at 0.659 (std=0.121). This pattern held in the second experiment ($\mathcal{D}$ishCover: 0.402, std=0.110; GPT-4o: 0.731, std=0.078).
@@ 7 Related Work
Our approach builds on recent parsing-based methods that guide LLMs to map natural language into structured forms, boosting LLM performance in tasks such as constituency parsing Tian et al. (2024a) and information extraction Zhao et al. (2023); Li et al. (2024). Similar to research integrating LLMs with Knowledge Graphs (KGs) to improve inference and reasoning Wang et al. (2024b); Feng et al. (2023); Jiang et al. (2023); Sun et al. (2023), we incorporate domain-specific knowledge into structured representations to provide LLMs with clearer, context-rich signals. Our work also aligns with models that parse text into structured knowledge, manipulate these representations, and (optionally) convert them into natural language to enhance LLM capabilities Yang et al. (2023); Besta et al. (2024); Zelikman et al. (2023); Zhang et al. (2025). We surprisingly show that similar techniques can augment creativity and diversity.
@@
Beyond improving LLMs, researchers have also explored their use as creative aids for writers Mirowski et al. (2023); Yuan et al. (2022); Chakrabarty et al. (2023); Wan et al. (2024), visual artists Ko et al. (2023), and even humorists Wu et al. (2025). However, while these tools boost users’ sense of creativity, their limited diversity may homogenize the ideas produced across different individuals Anderson et al. (2024).
@@
Recent work has explored the use of LLMs in the culinary domain Ma et al. (2024); Zhou et al. (2024); Ataguba and Orji (2025); Hwang et al. (2023a), including several early efforts in recipe generation H. Lee et al. (2020); Antô et al. (2020); Bień et al. (2020). Prior to the advent of LLMs, computational recipe generation focused primarily on proposing novel ingredient combinations, often neglecting complete cooking instruction generation Morris et al. (2012); Amorim et al. (2017); Cromwell et al. (2015); Varshney et al. (2019).
8 Discussion and Conclusions
We introduce a novel paradigm for enhancing creative generation in LLMs using structured recombination at the idea level rather than token-level.
Implications for AI Creativity and Structured Knowledge. A core insight from our work is that structured representations provide a mechanism to introduce meaningful variability at a higher level of abstraction. Attempts to enhance creativity in LLMs often focus on increasing token-randomness (e.g., through temperature), which is not sufficient for creativity Peeperkorn et al. (2024). We focus on recombination, a well-established principle in creativity research. By operationalizing this principle computationally, we show that incorporating structured knowledge enables a model to navigate the space of ideas more effectively, leading to novel and valuable outputs.
@@ 8 Discussion and Conclusions
Implications for Sampling. Another promising application of our approach is in structured sampling and search-based generation tasks. Sampling methods in LLMs often rely on stochastic techniques that operate at the token level (e.g., nucleus or top-k sampling) to introduce variability. In contrast, we diversify by sampling over abstract representations rather than raw tokens. We introduce a controlled yet flexible way of navigating idea spaces. This approach to sampling can enhance a range of applications that require balancing novelty and coherence.
Generalizability to Other Domains. Beyond cooking, our framework can generalize to any domain where structured knowledge can guide creative recombination, such as product design, scientific discovery, or narrative construction. Tree representations apply naturally to many of these domains, and our approach can also be extended to other structures, such as sequences, general graphs, or other formal structures. In addition to idea generation, our framework can support synthetic data simulation and augmentation, addressing data scarcity — a persistent bottleneck in training AI models for low-resource domains. In this setting, novelty scoring ensures varied outputs, while domain-specific constraints help maintain data validity.
Ultimately, our approach represents a step toward AI systems that do not just generate text but truly engage in creative synthesis. We hope to inspire further research in structured creativity and AI, highlighting the potential of structured recombination in augmenting human creativity and expanding the frontiers of AI-assisted innovation.
@@ Appendix A Fine-Tuned Sentence-BERT Model
To better handle recipe similarity, we fine-tuned a Sentence-BERT model555all-distilroberta-v1. . Our fine-tuning dataset consisted of 30K pairs of recipes with their new similarity scores, which equally weighted the original Sentence-BERT score and a Ruzicka-based similarity of the two recipes’ ingredient lists as computed by Mizrahi and Shahaf (2021). The dataset was divided into three equal subsets: (1) 10K pairs of recipes, each pair representing instances of the same dish, (2) 10K pairs of recipes from different dishes within the same category (e.g., carrot cake and cheesecake), and (3) 10K pairs of recipes from entirely different categories (e.g., a dessert and a salad).

### v2

@@ 6.1 Experimental Details
Here we describe the process of generating recipes for the GPT-4o baseline as well as the human evaluation setup.
Recipe Generation for the GPT-4o Baseline. LLMs are known to be sensitive to prompt paraphrases Sclar et al. (2023); Mizrahi et al. (2024); Voronov et al. (2024). To ensure a fair and competitive baseline, we conducted a thorough prompt design process. First, we constructed large, diverse pools of prompt paraphrases tailored for each experimental task. Specifically, we developed 104 prompts for Experiment 1 and 114 prompts for Experiment 2. These prompts varied systematically along several axes: structure (explicit steps vs. open-ended, explicit chef role vs. no role, etc.), length (concise vs. detailed instructions), creativity-oriented language, and creativity-related constraints. [^5] All prompt variations are publicly available alongside the dataset and code for full reproducibility.
We tested different temperature settings (ranging from 0.0 to 2.0) for a sample of these prompts. For $t>1$, GPT-4o often produced gibberish. In line with the findings of Peeperkorn et al. (2024), we observed improvements in novelty at higher temperatures, although they were rather subtle. We therefore selected $t=1$ as the highest temperature that produced coherent recipes.
To assess the diverse prompt set, we conducted a small evaluation. For Experiment 1, we generated a recipe for each of three sampled dish pairs and every prompt (resulting in 312 recipes). For Experiment 2, we generated three recipes independently per prompt (resulting in 342 recipes).
Qualitative analysis of the generated recipes revealed significant repetition: For experiment 1, GPT-4o consistently generated recipes that prepared each component separately and combined them at the end of the process in the same superficial manner. For example, all burger-waffle recipes involved cooking the burgers and waffles independently, then simply stacking the burger between two waffles. This observation is supported by high average Self-BLEU scores (BLEU-2=0.929, BLEU-3=0.877, BLEU-4=0.818), and high cosine similarity between recipes of the same pair (average $\sim$0.90 across sampled dish pairs).
For experiment 2, GPT-4o frequently reused known culinary dish concepts with minor variations (e.g., marinated proteins served with quinoa/rice in approximately 60 recipes). This observation is again supported by high self-BLEU scores (BLEU-2=0.912, BLEU-3=0.829, BLEU-4=0.739), and high similarity to the large corpus of existing recipes, with average cosine similarity to nearest corpus neighbor of 0.853 (std=0.035).
Due to this pronounced repetition, using multiple prompts in the experiments risked redundant outputs, weakening the baseline. Therefore, two culinary experts reviewed the GPT-4o outputs to select the single best-performing prompt per experiment – the one that consistently produced the most creative and varied recipes. Importantly, this means that the GPT-4o baseline was explicitly optimized for the experiments’ evaluation criterion (creativity as judged by human experts), thus giving it a potential advantage over $\mathcal{D}$ishCover.
Finally, we generated the baseline outputs for the experiments using these best-performing prompts within a single chat session, instructing GPT-4o to produce recipes differing from prior outputs to encourage diversity. See the chosen prompts for both experiments in Appendix G.
@@
Human Evaluation Setup. We used Prolific to recruit and manage a total of 48 participants for both experiments. We pre-screened participants based on their cooking experience and frequency, comfort with adjusting recipes, and ability to judge creative outcomes. Participants were paid an estimated hourly rate of £9, adhering to ethical compensation standards. The participants rated randomly selected recipes drawn from the two models on various aspects related to novelty and value (see the full list of questions in Appendix H). To reduce cognitive load, participants initially viewed concise recipe summaries (also generated by GPT-4o), with the option to examine complete recipes if desired. Each recipe was rated by five annotators. Final scores were determined using the median of annotators’ ratings, a standard approach for ordinal data to handle outliers. Novelty and value scores were calculated as the mean ratings of their respective questions; the value score was additionally binarized using a threshold of 4 (on a 1–5 scale) across three related questions. On average, participants rated 31.25 recipes (std=33.827). Participants who consistently completed ratings too quickly (<<45 seconds per recipe) were excluded to ensure data quality.
Human Evaluation Reliability. To confirm the robustness of our human evaluation results, given the inherent subjectivity involved in judging creativity, we computed two metrics commonly used in subjective annotation tasks: Interquartile Range (IQR), capturing numeric consistency among annotators, and Krippendorff’s alpha, assessing overall agreement reliability.
[…]
  • Krippendorff’s alpha: To further assess inter-rater reliability for ordinal data, we computed Krippendorff’s alpha, obtaining values of 0.201 (Experiment 1) and 0.209 (Experiment 2). While modest, these alpha values are consistent with expectations for subjective tasks like creativity evaluation, where individual interpretations naturally differ, even when numerical responses show overall consistency.
@@ 6.2 Diversity in Generated Recipes
This difference was also reflected in the average ingredient and word counts: GPT-4o’s recipes were nearly twice as long and complex as $\mathcal{D}$ishCover’s. In Exp. 1, $\mathcal{D}$ishCover had on average 12.3 ingredients (std=3.71) and 240.28 words (std=47.98), compared to GPT-4o’s 24.98 ingredients (std=5.2) and 465.84 words (std=64.8). In Exp. 2, $\mathcal{D}$ishCover used 13.32 ingredients (std=4.39) and 264.15 words (std=51.38) on average, compared to GPT-4o’s 23.79 ingredients (std=3.25) and 371.7 words (std=42.92). For reference, recipes in the Recipe1M+ corpus averaged only 9.33 ingredients (std=4.31) and 168.53 words (std=104.79). This suggests that GPT-4o tends to generate separate sub-dishes, each with its own ingredients and instructions, and then combine them into a final dish rather than blending the culinary ideas.
@@
GPT-4o showed a strong fixation on certain structures and ingredients. When asked to merge two dishes, it often followed the same approach in multiple attempts. For instance, when asked to fuse “lentil soup” and “jam”, GPT-4o repeatedly generated variations of lentil soup served alongside a separately prepared jam, plated in the same way (e.g., placing a dollop of jam in the center of the soup). This fixation persisted when asked to combine specific given recipes, and even when asked to merge broader dish types (e.g., a general soup with a general dip).
Similarly, when asked to suggest a general creative recipe, GPT-4o repeatedly used the same (uncommon) ingredients: 56% of its recipes included smoked paprika (compared to just 0.375% in the repository), 47% used mint, 41% maple syrup, and 29% pomegranate molasses. Figure 5 shows ingredient frequencies in recipes of GPT-4o and $\mathcal{D}$ishCover against the repository baseline. While $\mathcal{D}$ishCover’s distribution aligns closely with the original data, GPT-4o has a strong bias toward certain low-frequency ingredients.
Figure 5: Comparing ingredient frequencies in the original recipe repository (in pink) and the models’ generated recipes (Exp. 2). The black histogram shows $\mathcal{D}$ishCover closely follows the repository’s distribution. GPT-4o (blue) shows spikes, highlighting bias to certain ingredients.
@@
Quantifying Lexical Diversity via Self-BLEU.
We computed self-BLEU scores across the generated recipes for both experiments. GPT-4o exhibited significantly higher redundancy, reflected in higher self-BLEU scores: In Experiment 1, GPT-4o’s outputs scored 0.726 (BLEU-2), 0.655 (BLEU-3), and 0.599 (BLEU-4), while $\mathcal{D}$ishCover’s scores were substantially lower: 0.423, 0.281, and 0.192, respectively. Similarly, in Experiment 2 GPT-4o scored 0.902, 0.828, and 0.753, compared to $\mathcal{D}$ishCover’s 0.778, 0.619, and 0.475. The differences are more pronounced in Experiment 1 due to the smaller reference set (5 recipes per pair, vs. 100 recipes).
Quantifying Intra-Set Diversity.
We also analyzed the embeddings of the generated recipes using a fine-tuned Sentence-BERT model specialized for cooking recipes (see Appendix A). Figure 6 presents a t-SNE visualization for the first experiment, where each shape represents a dish pair, and color indicates the model ($\mathcal{D}$ishCover in pink, GPT-4o in blue). GPT-4o’s recipes formed tight clusters for each pair, whereas $\mathcal{D}$ishCover’s were more dispersed. A similar pattern emerges in experiment 2 (Figure 7). By construction, $\mathcal{D}$ishCover’s embeddings are widely distributed. The interesting thing to notice, however, is that GPT-4o’s embeddings tend to cluster closely. Figure 8 reinforces these findings with heatmaps of cosine similarities among all recipe embeddings. In the first experiment, $\mathcal{D}$ishCover’s outputs had an average similarity of 0.387 (std=0.120), while GPT-4o’s were higher at 0.659 (std=0.121). Results were similar in Experiment 2 ($\mathcal{D}$ishCover: 0.402, std=0.110; GPT-4o: 0.731, std=0.078).
Quantifying Diversity with Respect to an External Corpus.
We also assessed the similarity of generated recipes to known recipes by computing cosine similarity between a recipe’s embedding and that of its nearest neighbor in the Recipe1M+ corpus. $\mathcal{D}$ishCover’s recipes consistently exhibited greater distance from existing recipes, with an average similarity of 0.774 (std=0.061), significantly lower than GPT-4o’s average of 0.851 (std=0.027). Additionally, 91% of $\mathcal{D}$ishCover’s recipes had similarity below 0.85, compared to only 47% of GPT-4o’s. Notably, 38% of $\mathcal{D}$ishCover’s outputs had similarity score below 0.75, while none of GPT-4o’s did.
Figure 6: t-SNE visualization for recipe embeddings, experiment 1. Shapes denote dish pairs, colors and fill style denote models. The filled clusters (GPT-4o) of the same shape appear tight and localized, while the unfilled markers ($\mathcal{D}$ishCover) are scattered, showing higher diversity across $\mathcal{D}$ishCover’s outputs.
Figure 7: t-SNE visualization for recipe embeddings, experiment 2. The x markers represent 1K random recipes from the general repository. Again, $\mathcal{D}$ishCover’s outputs are more diverse.
@@ 7 Related Work
Our approach builds on recent parsing-based methods that guide LLMs to map natural language into structured forms, boosting LLM performance in tasks such as constituency parsing Tian et al. (2024a) and information extraction Zhao et al. (2023); Li et al. (2024). Similar to research integrating LLMs with Knowledge Graphs (KGs) to improve inference and reasoning Feng et al. (2023); Jiang et al. (2023); Sun et al. (2023); Wang et al. (2024b), we incorporate domain-specific knowledge into structured representations to provide LLMs with clearer, context-rich signals.
Our work also aligns with models that parse text into structured knowledge, manipulate these representations, and (optionally) convert them into natural language to enhance LLM capabilities Yang et al. (2023); Zelikman et al. (2023); Besta et al. (2024); Zhang et al. (2025). We surprisingly show that similar techniques can augment creativity and diversity. In parallel, the structured representations we use share a conceptual similarity with latent variable models, which learn to encode inputs into structured latent forms and apply transformations to generate viable alternatives Kusner et al. (2017); Dai et al. (2018); Zhang et al. (2019). Unlike these approaches, our method does not learn a latent space and requires no training.
@@
Common computational approaches for recombining ideas include conceptual blending frameworks, where merging is guided by heuristics or rules Fauconnier and Turner (2003); Pereira and Cardoso (2006); Veale and Cardoso (2019); genetic algorithms, which use random crossover points Corne and Bentley (2001); Cho (2002); Dennis and Stella (2011); and latent space blending, which interpolates between learned representations but lacks explicit structure Sarkar and Cooper (2021); Yee-King (2022); Zhou et al. (2025). Our approach differs by recombining structured representations through an edit distance algorithm, thereby providing explicit and controllable blending.
@@
Beyond improving LLMs, researchers have also explored their use as creative aids for writers Yuan et al. (2022); Mirowski et al. (2023); Chakrabarty et al. (2024b); Wan et al. (2024), visual artists Ko et al. (2023), and even humorists Wu et al. (2025). However, while these tools boost users’ sense of creativity, their limited diversity may homogenize the ideas produced across different individuals Anderson et al. (2024).
@@
Recent work has explored the use of LLMs in the culinary domain Hwang et al. (2023a); Ma et al. (2024); Zhou et al. (2024); Ataguba and Orji (2025); Thomas et al. (2025), including several early efforts in recipe generation H. Lee et al. (2020); Antônio et al. (2020); Bień et al. (2020). Prior to the advent of LLMs, computational recipe generation focused primarily on proposing novel ingredient combinations, often neglecting complete cooking instruction generation Morris et al. (2012); Cromwell et al. (2015); Amorim et al. (2017); Varshney et al. (2019).
8 Discussion & Future Work
Generalizability to Other Domains. Extending our approach beyond the culinary domain presents important considerations. Our pipeline fundamentally relies on (1) meaningful structured representations, and (2) the ability to define or approximate reliable criteria for assessing value.
As discussed in Section 4, procedural texts naturally exhibit tree-like structures: leaves often correspond to objects, and internal nodes represent operations performed on them. In storytelling, graphs are commonly used to model narrative arcs and plot dynamics Elson (2012); Valls-Vargas et al. (2017). In domains such as drug discovery, molecular structures are inherently graph-based, with atoms as nodes and bonds as edges Ivanciuc and Balaban (2000). Music also allows structured, hierarchical representations Good (2001); Cuthbert and Ariza (2010). Other domains, such as poetry or product ideation, lack obvious or standardized structured representations, making the applicability of our technique less immediate.
Value assessment poses another challenge. In recipes, we leverage the FlavorDB dataset as a proxy metric for taste. Our method can be adapted to other domains that employ similar domain-specific metrics. For example, in music, one can predict how humans perceive a combination of sounds (based on psychoacoustic consonance heuristics). Domains related to physical construction (furniture, architecture) could apply automated feasibility checks through CAD software. For programming and game design, functionality could similarly be verified automatically through simulation. In domains where value assessments are inherently subjective and difficult to automate, domain expert evaluation could be integrated into the pipeline, enabling hybrid approaches.
Structured Recombination in LLMs: Implications for AI Creativity and Sampling. A core insight of our work is that structured recombination provides a controllable mechanism to introduce meaningful variability at a higher level of abstraction, rather than the lexical, token level.
To date, attempts to enhance creativity in LLMs often focused on increasing token-randomness (e.g., via temperature), which is not sufficient for creativity Peeperkorn et al. (2024). Structured recombination enables a model to explore the creative space of ideas more deliberately and effectively. Together with modules for assessing value and novelty this can lead to more creative outputs.
@@ 8 Discussion & Future Work
Another promising application of structured recombination is in structured sampling and search-based generation tasks. Standard sampling methods in LLMs Fan et al. (2018); Holtzman et al. (2019) often rely on stochastic techniques that operate at the token level (e.g., nucleus or top-k sampling) to introduce variability. In contrast, we diversify by sampling over abstract representations rather than raw tokens. We introduce a controlled yet flexible way of navigating idea spaces, which can enhance a range of applications that require deliberate exploration of diverse concepts.
9 Conclusions
In this work, we addressed the persistent challenge of generating creative, diverse outputs. We proposed a novel approach that leverages structured representations to enhance creativity. Rather than relying on superficial token-level variation, we perform cognitively-inspired manipulations – specifically, recombining structured representations of existing concepts. We demonstrated our paradigm in the culinary domain through $\mathcal{D}$ishCover, a model designed to generate creative and diverse recipes, and empirically showed significant improvements over GPT-4o in terms of creativity. Ultimately, our approach represents a step toward pushing LLMs beyond surface-level variation, opening the path to richer and more controllable creative generation. We hope this work inspires further research into structured creativity in AI, and we invite the community to build upon this paradigm across diverse domains.
Acknowledgements
We thank the reviewers and action editor for their insightful comments. We further thank Dana Aviran and Eitan Stern for developing the Sentence-BERT fine-tuned model on recipe data, and to the members of Hyadata Lab and SLAB at the Hebrew University of Jerusalem for their thoughtful remarks. This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant no. 852686, SIAM), by the Israeli Ministry of Science and Technology (grant no. 7256), and by the Koret Foundation grant for Smart Cities and Digital Living.
@@ Appendix A Fine-Tuned Sentence-BERT Model
To better handle recipe similarity, we fine-tuned a Sentence-BERT model.666all-distilroberta-v1. Our fine-tuning dataset consisted of 30K pairs of recipes with their new similarity scores, which equally weighted the original Sentence-BERT score and a Ruzicka-based similarity of the two recipes’ ingredient lists as computed by Mizrahi and Shahaf (2021). The dataset was divided into three equal subsets: (1) 10K pairs of recipes, each pair representing instances of the same dish, (2) 10K pairs of recipes from different dishes within the same category (e.g., carrot cake and cheesecake), and (3) 10K pairs of recipes from entirely different categories (e.g., a dessert and a salad).